24 January 2016

Mtn Manager: Mike Wagner
OSA: D. Gonzalez Huerta
InstSp: J. Morris
Observers: B. Rothberg, D. Thompson, A. Cardwell, S. Walsh, O. Kuhn (remote)

Night started with winds at the limit and clouds. Did not open. D. Thompson encountered a LUCI2 mask problem - no fault, but the mask simply would not move. J. Reidl called and working on it. There were flexure compensation checks that could have been run closed dome. Instead, with dome closed, running the final set of MODS2 cals needed before the MODS2 blue fiber can be switched back from M2.BC to the PCI computer.

Rick needed some MODS2 calibrations for the pipeline. Taking these now at EL=60 deg and PA=0 (RDG rotator at angle 195 deg).
  • pinholeSlit.cal
    • r 2-8 and b 2-7
  • sieveSnap.pro:
    • r=9 and b=8
  • sudokuMask:
    • r=10-16 b=9-14

06:00 UT: Maybe clearing a bit...

-- %USERSIG{OlgaKuhn - 2016-01-24}%

Comments

 
Topic revision: r2 - 24 Jan 2016, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback